OpenStudy (anonymous):

How do I do this problem? n an 1 1 2 6 3 11 4 16 When n equals 8, what does an equal?

OpenStudy (anonymous):

You can tell it's an arithmetic sequence because \[ \large a_{n+1} -a_n=5 \]

OpenStudy (anonymous):

Arithmetic sequences have the form: \[ a_n = a_1 + (n-1)d \]

OpenStudy (anonymous):

In this case \(d = 5\) and \(a_1=1\), so our equation is \[ a_n = 1 + (n-1)5 \]

Parth (parthkohli):

But that is not the question... your formula is:\[a_n = a_1 + (n - 1)d\]The \(d\) was stated by @wio: \(5\).

OpenStudy (anonymous):

So you want to substitute \(n=8\) into the equation to find \(a_8\).

Parth (parthkohli):

So, the formula is:\[a_n = 1 + 5(n - 1)\]Now substitute \(n = 8\).

OpenStudy (anonymous):

So the answer would be 36, correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

:D Thank you both, yall were A LOT of help!! <3